# AI for Science

### v1

@@ 1 Introduction
Peptide-based therapeutics are gaining momentum in modern drug discovery due to their high specificity, tunable bioactivity, and ability to target challenging protein–protein interactions[1]. Cyclic peptides, in particular, offer enhanced stability and binding affinity from their constrained conformations and the incorporation of non-natural amino acids (NNAAs)[2]. Despite this promise, designing peptides with multiple desired properties, such as solubility and binding affinity, remains a formidable challenge due to the vastness and complexity of the design space and the scarcity of labeled data for optimization[3].
The application of artificial intelligence (AI), especially generative models, has shown remarkable progress in addressing these challenges[4, 5, 6, 7]. Tools such as PepTune[8] and PepINVENT[9] represent two notable advances in the field. PepTune leverages discrete masked diffusion models and Monte Carlo Tree Search (MCTS) for multi-objective optimization of peptide SMILES strings, enabling the design of chemically modified and cyclic peptides in a discrete representation space. Similarly, PepINVENT extends the REINVENT[10, 11] platform to peptides with a masked language model as an idea generator, enabling peptide optimization at specified monomer position. While the impact of these tools on real-world projects remains to be seen, a persistent limitation of current peptide design frameworks is their lack of interpretable reasoning during generation. Without insights into why a particular sequence of modifications enhances a given design objective, such models can be difficult to trust or adapt to program-specific constraints.
Inspired by recent advancement in reasoning in Large Language Models (LLMs)[12, 13], we propose a novel framework that integrates LLM with explicit chain-of-thought (CoT) supervised fine-tuning (SFT) and reinforcement learning (RL) to improve interpretable property controlled cyclic peptide optimization. Unlike previous work, our approach directly leverages monomer-level reasoning during sequence generation, offering improved interpretability and modularity, which is critical for iterative design in real-world therapeutic applications. Specifically, our contributions include:
@@ 2 Related Work
Designing functional peptides, including cyclic peptides, is challenging due to the immense sequence space and complex property requirements. Recent work has explored large language models (LLMs), variational autoencoders, and other deep generative models as the generator for new peptides, and using techniques such as RL[9], MCTS[8, 4], genetic algorithms [14] and LLM[5] to guide exploration of peptide sequence space toward desired properties.
@@
The majority of work represents peptide as linear sequence in HELM[15] or SMILES[16], and apply generative modeling to design peptides with tailored therapeutic properties. Several models, such as HELM-GPT[7] and AMP-Designer[6], are GPT-based models that operate at monomer level using HELM representation. SMILES-based representation is adopted by PepINVENT and PepTune, both leverage masked language models for diverse peptide generation. Some work utilizes the conversion between different representation forms to enable generation in sequence while evaluation with three-dimensional structural information[4, 17]. More related work can be found in reviews by Wan, Zhai, et al. [18, 19, 20].
@@
Recent work[11, 21, 22, 23] has shown that reinforcement learning (RL) is an effective strategy for multi-property optimization in therapeutic entity design, enabling goal-directed discovery in multiple modalities including small molecule, peptide, and protein. Many works[6, 8, 9] integrate RL modules to fine-tune generative models with feedback from property predictors to optimize traits like cell permeability, antimicrobial activity, and binding affinity. Techniques range from policy gradient and reward shaping to inference-time search strategies like Monte Carlo Tree Guidance[8], allowing models to balance trade-offs across multiple, often conflicting therapeutic objectives.
@@
Chain-of-Thought (CoT)[13] aims to enhance the reasoning capability of LLMs by generating intermediate reasoning steps before arriving at a final answer and is now a key technology in enhancing reasoning capability in modern LLMs[12, 24]. A few recent works have extended CoT reasoning to address biochemistry related problems, including chemical question-answering tasks[25], protein-protein interaction prediction[26], molecular structural understanding[27], property prediction[28]. Directly incorporating CoT into SFT of LLMs for chemical optimization remains unseen to the best of our knowledge.
@@ 3 Methodology
The overview of our approach is depicted in Figure.1. We first augmented the raw peptide data from CycPeptMPDB database [29] through single position random mutation of peptide HELM [15] code to generate peptide pairs with single monomer difference, and then used internal QSAR model [30] to get property values of all peptides. High quality peptide pairs were used for constructing training data for CoT SFT of pretrained LLM [31]. To further improve property control of generated peptides, an RL strategy was designed to further optimize the model. Details of our training data construction strategy, design of CoT prompt, and RL approach are introduced below.
@@ 3.1 Preparing synthetic training dataset
We constructed the dataset starting from 5,530 head-to-tail cyclic peptide sequences extracted from the CycPeptMPDB database [29]. After removing sequences that can not be processed by CycloPs [32], 3,778 peptides composed solely of natural $\alpha$-amino-acid (AA) monomers were retained. To expand the chemical space, we incorporated non-natural $\alpha$-amino acids (NNAAs) from the virtual library developed by Amarasinghe et al. [33]. This reaction-based library contains approximately 380,000 synthesizable NNAAs, from which we used a representative subset of 10,000. Combined with 385 natural monomers from CycPeptMPDB, this resulted in a building block vocabulary of 10,385 unique monomers.
@@
Peptide augmentation was carried out using CycloPs [32]. For each of the 3,778 natural peptides, we applied 100 independent point mutations at randomly selected non-terminal positions in their HELM sequences [15]. In each mutation, the original residue was replaced by a randomly sampled monomer from the monomer vocabulary. For every resulting sequence, we recorded the mutated peptide, the mutation index, and the pre- and post-mutation monomer identities. This procedure yielded a raw dataset of approximately 380,000 unique cyclic peptides, and each mutated peptide and its original counterpart naturally became a peptide pair with only single monomer difference.
@@
All peptides were annotated using a multi-task message-passing QSAR model built upon Chemprop [30], trained for ADMET-relevant property prediction on internal data as reported by Chen et al. [34]. We focused on three endpoints predicted by QSAR: (1) LogD (LogD_HPLC_pH_7) representing lipophilicity at pH 7; (2) MRT (Rat_MRT(hrs)) representing mean residence time in rats, reported in hours; and (3) SIF (SIF_halflife_10mg_per_ml(hrs)) representing the peptide half-life in simulated intestinal fluid.
@@ 3.3 Reinforcement learning with pharmacology-aware rewards
For the RL training phase, We fine-tuned our model $\pi_{\theta}$ using Group Relative Policy Optimization (GRPO) [35] to enhance the reasoning process. Following the CoT prompt template, each training sample $x\in\mathcal{D}$ ($\mathcal{D}$ represents training dataset) consists of a SMILES string to be edited $x_{s}$ and a textual prompt $x_{t}$. Given input $x=(x_{s},x_{t})$, the model samples a group of candidate outputs $\{o^{(1)},\ldots,o^{(G)}\}\sim\pi_{\theta}(y\mid x)$, where each $o^{(i)}$ includes both reasoning trace and final answer. We define a customized reward function to score each output. Relative scores $\{R_{i}\}$ are then computed and normalized to obtain advantages $\{A_{i}\}$, which are used to update $\pi_{\theta}$ following the GRPO objective (Eq. (1)) as follows:
@@
To guide reinforcement learning toward the generation of pharmacologically promising molecules, we designed a reward function that balances property desirability, structural similarity, and diversity control by adapted partially from REINVENT[11]. The final reward $R$ for a candidate peptide is computed as:
@@
Property desirability. We employed a Chemprop-based multi-task QSAR model[30] to predict three key ADMET properties: LogD, MRTRat, and SIF $t_{1/2}$. Instead of using raw regression outputs, we applied a smooth sigmoid-based transformation to capture whether a prediction exceeds a predefined threshold:
@@
Validity measures the fraction of generated sequences that are chemically valid by RDKit[36] and successfully predicted by the QSAR property predictor. A higher score indicates fewer invalid or corrupted molecules. Novelty quantifies the percentage of uniquely generated molecules that are not present in the training dataset, capturing the model’s ability to generate novel compounds. Uniqueness reflects the ratio of distinct molecules among all generated samples, measuring diversity and redundancy.
@@ 4.1 CoT SFT and RL empowered PepThink-R1 to be an effective peptide optimizer
This experiment shows that PepThink-R1 is very effective in optimizing peptides towards multi-property improvement. Its capability mainly comes from the CoT SFT and RL steps in our workflow. Considering that lab data rarely shows LogD above 6 and SIF above 30hr, it is important to not take the predicted property value as they are, but instead, should interpret the results as relative improvement over benchmarks under same QSAR model. Performance of internal QSAR model on the data set used in this work could be impacted by many factors, such as domain applicability and activity cliff [37, 38]. Hence, this work intrinsically is also limited by those factors given that it was built upon QSAR model. Our exploration of RL is not comprehensive, and we observed that more gains could be achieved through further refining RL reward model.
@@ 4.2 Our models outperform general LLMs
To evaluate PepThink-R1, we considered a series of general-purpose LLMs including both reasoning and non-reasoning models for comparison. Specifically, we include GPT-4o [39], DeepSeek-R1-Llama-8B [40], Qwen3-30B-Thinking-2507 [41] and GPT-5 [42] as LLM baselines. The two GPT models are accessed via API and all other baselines are accessed through open-source weights. The API version of GPT-4o is set to ’gpt-4o-2024-08-06’ and the reasoning effort of GPT-5 is set to minimal. For benchmarking, we adopted vllm[43] as the inference framework. We used a consistent set of inference hyperparameters. The sampling temperature is set to 0.95, the top-p is set to 0.7 and we restrict the maximum number of generated tokens to 4096.
@@ 4.3 Case study on PepThink-R1 versus PepINVENT
We included PepINVENT, a transformer-based model as a domain-specific baseline and conducted a small case study with four seed peptides to qualitatively compare PepINVENT and PepThink-R1 under challenging settings. Another notably model to compare with is PepTune [8], but we were not able to use it due to limited authorization to for-profit commercial entities. For each seed peptide, we first ranked PepINVENT candidates by the composite total_score derived from LogD, MRT, and SIF, and retained the top-scoring candidate as the PepINVENT examples. For PepThink-R1, we selected one candidate per seed that met the “high” thresholds on all three properties (LogD $\geq$ 4.2, MRT $\geq$ 1.63 h, SIF $t_{1/2}$ $\geq$ 10.1 h). We also include the original seed peptide as the baseline.
@@ 4.4 Interpretable reasoning process by PepThink-R1
Table. 3 shows the exact text generated by PepThink-R1 with visualization of chemical structures of monomers involved in each case. For peptide 1, it mentions replacing glycine with [NH2+][C@@H](Cc1cc(C(C)(C)C)cc(c1Cl)Cl)C(=O)[O-]. The leaving monomer has a matched residue in peptide 1 and the matched residue is replaced by residue of the proposed monomer in peptide generated by PepThink-R1. For peptide 2, the reasoning output instructs replacing leucine with the new monomer which is strictly reflected in the structural change. In both cases, there are more monomer changes proposed by PepThink-R1 that are not captured in the reasoning text. PepINVENT does not have reasoning process, hence, we turned to general reasoning LLMs - GPT-4o and GPT-5 - for comparison. Reasoning outputs from two GPT models for peptide 1 are captured in Table 5. Both outputs from GPT-4o and GPT-5 are generally reasonable high-level strategies for optimizing peptides towards better LogD, MRT, and SIF, and these two outputs share many commonalities, including proposing N-methylation and introducing tert-butyl group. They also roughly cover strategies adopted by PepINVENT and PepThink-R1, but missing detailed monomer structures captured in PepThink-R1 reasoning text. One caveat with GPT reasoning outputs is that they both lack the specific details with the given peptide, which illustrates a limitation of general purpose LLMs and may explain their poor performance (Table.2). The self-consistency discrepancy between great GPT strategy in reasoning output and poor quantitative metrics of resulting peptide SMILES shown in Figure 7 could indicate that either the reasoning process does not reflect the real thinking process of the GPT models, or GPT models are not trained to faithfully carry out their own strategies[44, 45]. Conversely, PepThink-R1 reasoning output is interpretable, faithfully rooted in the specific case, and also accurately reflected in the final peptide SMILES. Comparing to GPT models, PepThink-R1’s limitation is that its reasoning fails to capture all changes and does not reflect as broad a strategy as GPT is capable of.

### v2

@@ 1 Introduction
Peptide-based therapeutics are gaining momentum in modern drug discovery due to their high specificity, tunable bioactivity, and ability to target challenging protein–protein interactions[wang2022therapeutic]. Cyclic peptides, in particular, offer enhanced stability and binding affinity from their constrained conformations and the incorporation of non-natural amino acids (NNAAs)[ji2024cyclic]. Despite this promise, designing peptides with multiple desired properties, such as solubility and binding affinity, remains a formidable challenge due to the vastness and complexity of the design space and the scarcity of labeled data for optimization[muttenthaler2021trends].
The application of artificial intelligence (AI), especially generative models, has shown remarkable progress in addressing these challenges[lin2025highplay, gong2025pepseek, wang2025discovery, xu2024helm]. Tools such as PepTune[tang2025peptune] and PepINVENT[geylan2025pepinvent] represent two notable advances in the field. PepTune leverages discrete masked diffusion models and Monte Carlo Tree Search (MCTS) for multi-objective optimization of peptide SMILES strings, enabling the design of chemically modified and cyclic peptides in a discrete representation space. Similarly, PepINVENT extends the REINVENT[olivecrona2017molecular, loeffler2024reinvent] platform to peptides with a masked language model as an idea generator, enabling peptide optimization at specified monomer position. While the impact of these tools on real-world projects remains to be seen, a persistent limitation of current peptide design frameworks is their lack of interpretable reasoning during generation. Without insights into why a particular sequence of modifications enhances a given design objective, such models can be difficult to trust or adapt to program-specific constraints.
Inspired by recent advancement in reasoning in Large Language Models (LLMs)[guo2025deepseek, wei2022chain], we propose a novel framework that integrates LLM with explicit chain-of-thought (CoT) supervised fine-tuning (SFT) and reinforcement learning (RL) to improve interpretable property controlled cyclic peptide optimization. Unlike previous work, our approach directly leverages monomer-level reasoning during sequence generation, offering improved interpretability and modularity, which is critical for iterative design in real-world therapeutic applications. Specifically, our contributions include:
@@ 2 Related Work
Designing functional peptides, including cyclic peptides, is challenging due to the immense sequence space and complex property requirements. Recent work has explored large language models (LLMs), variational autoencoders, and other deep generative models as the generator for new peptides, and using techniques such as RL[geylan2025pepinvent], MCTS[tang2025peptune, lin2025highplay], genetic algorithms [krishnan2025generative] and LLM[gong2025pepseek] to guide exploration of peptide sequence space toward desired properties.
@@
The majority of work represents peptide as linear sequence in HELM[zhang2012helm] or SMILES[weininger1988smiles], and apply generative modeling to design peptides with tailored therapeutic properties. Several models, such as HELM-GPT[xu2024helm] and AMP-Designer[wang2025discovery], are GPT-based models that operate at monomer level using HELM representation. SMILES-based representation is adopted by PepINVENT and PepTune, both leverage masked language models for diverse peptide generation. Some work utilizes the conversion between different representation forms to enable generation in sequence while evaluation with three-dimensional structural information[lin2025highplay, chen2025target]. More related work can be found in reviews by Wan, Zhai, et al. [wan2022deep, zhai2025artificial, wan2024machine].
@@
Recent work[loeffler2024reinvent, wang2024reinforcement, wang2023self, popova2018deep] has shown that reinforcement learning (RL) is an effective strategy for multi-property optimization in therapeutic entity design, enabling goal-directed discovery in multiple modalities including small molecule, peptide, and protein. Many works[wang2025discovery, tang2025peptune, geylan2025pepinvent] integrate RL modules to fine-tune generative models with feedback from property predictors to optimize traits like cell permeability, antimicrobial activity, and binding affinity. Techniques range from policy gradient and reward shaping to inference-time search strategies like Monte Carlo Tree Guidance[tang2025peptune], allowing models to balance trade-offs across multiple, often conflicting therapeutic objectives.
@@
Chain-of-Thought (CoT)[wei2022chain] aims to enhance the reasoning capability of LLMs by generating intermediate reasoning steps before arriving at a final answer and is now a key technology in enhancing reasoning capability in modern LLMs[guo2025deepseek, yeo2025demystifying]. A few recent works have extended CoT reasoning to address biochemistry related problems, including chemical question-answering tasks[ouyang2023structured], protein-protein interaction prediction[jin2024prollm], molecular structural understanding[jang2024chain], property prediction[jin2025effective]. Directly incorporating CoT into SFT of LLMs for chemical optimization remains unseen to the best of our knowledge.
@@ 3 Methodology
The overview of our approach is depicted in Figure.1. We first augmented the raw peptide data from CycPeptMPDB database [li2023cycpeptmpdb] through single position random mutation of peptide HELM [zhang2012helm] code to generate peptide pairs with single monomer difference, and then used internal QSAR model [heid2023chemprop] to get property values of all peptides. High quality peptide pairs were used for constructing training data for CoT SFT of pretrained LLM [yu2024llasmol]. To further improve property control of generated peptides, an RL strategy was designed to further optimize the model. Details of our training data construction strategy, design of CoT prompt, and RL approach are introduced below.
@@ 3.1 Preparing synthetic training dataset
We constructed the dataset starting from 5,530 head-to-tail cyclic peptide sequences extracted from the CycPeptMPDB database [li2023cycpeptmpdb]. After removing sequences that can not be processed by CycloPs [duffy2011cyclops], 3,778 peptides composed solely of natural $\alpha$-amino-acid (AA) monomers were retained. To expand the chemical space, we incorporated non-natural $\alpha$-amino acids (NNAAs) from the virtual library developed by Amarasinghe et al. [amarasinghe2022virtual]. This reaction-based library contains approximately 380,000 synthesizable NNAAs, from which we used a representative subset of 10,000. Combined with 385 natural monomers from CycPeptMPDB, this resulted in a building block vocabulary of 10,385 unique monomers.
@@
Peptide augmentation was carried out using CycloPs [duffy2011cyclops]. For each of the 3,778 natural peptides, we applied 100 independent point mutations at randomly selected non-terminal positions in their HELM sequences [zhang2012helm]. In each mutation, the original residue was replaced by a randomly sampled monomer from the monomer vocabulary. For every resulting sequence, we recorded the mutated peptide, the mutation index, and the pre- and post-mutation monomer identities. This procedure yielded a raw dataset of approximately 380,000 unique cyclic peptides, and each mutated peptide and its original counterpart naturally became a peptide pair with only single monomer difference.
@@
All peptides were annotated using a multi-task message-passing QSAR model built upon Chemprop [heid2023chemprop], trained for ADMET-relevant property prediction on internal data as reported by Chen et al. [chen2025data]. We focused on three endpoints predicted by QSAR: (1) LogD (LogD_HPLC_pH_7) representing lipophilicity at pH 7; (2) MRT (Rat_MRT(hrs)) representing mean residence time in rats, reported in hours; and (3) SIF (SIF_halflife_10mg_per_ml(hrs)) representing the peptide half-life in simulated intestinal fluid.
@@ 3.3 Reinforcement learning with pharmacology-aware rewards
For the RL training phase, We fine-tuned our model $\pi_{\theta}$ using Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical] to enhance the reasoning process. Following the CoT prompt template, each training sample $x\in\mathcal{D}$ ($\mathcal{D}$ represents training dataset) consists of a SMILES string to be edited $x_{s}$ and a textual prompt $x_{t}$. Given input $x=(x_{s},x_{t})$, the model samples a group of candidate outputs $\{o^{(1)},\ldots,o^{(G)}\}\sim\pi_{\theta}(y\mid x)$, where each $o^{(i)}$ includes both reasoning trace and final answer. We define a customized reward function to score each output. Relative scores $\{R_{i}\}$ are then computed and normalized to obtain advantages $\{A_{i}\}$, which are used to update $\pi_{\theta}$ following the GRPO objective (Eq. (1)) as follows:
@@
To guide reinforcement learning toward the generation of pharmacologically promising molecules, we designed a reward function that balances property desirability, structural similarity, and diversity control by adapted partially from REINVENT[loeffler2024reinvent]. The final reward $R$ for a candidate peptide is computed as:
@@
Property desirability. We employed a Chemprop-based multi-task QSAR model[heid2023chemprop] to predict three key ADMET properties: LogD, MRTRat, and SIF $t_{1/2}$. Instead of using raw regression outputs, we applied a smooth sigmoid-based transformation to capture whether a prediction exceeds a predefined threshold:
@@
Validity measures the fraction of generated sequences that are chemically valid by RDKit[landrum2013rdkit] and successfully predicted by the QSAR property predictor. A higher score indicates fewer invalid or corrupted molecules. Novelty quantifies the percentage of uniquely generated molecules that are not present in the training dataset, capturing the model’s ability to generate novel compounds. Uniqueness reflects the ratio of distinct molecules among all generated samples, measuring diversity and redundancy.
@@ 4.1 CoT SFT and RL empowered PepThink-R1 to be an effective peptide optimizer
This experiment shows that PepThink-R1 is very effective in optimizing peptides towards multi-property improvement. Its capability mainly comes from the CoT SFT and RL steps in our workflow. Considering that lab data rarely shows LogD above 6 and SIF above 30hr, it is important to not take the predicted property value as they are, but instead, should interpret the results as relative improvement over benchmarks under same QSAR model. Performance of internal QSAR model on the data set used in this work could be impacted by many factors, such as domain applicability and activity cliff [sheridan2015relative, sheridan2020experimental]. Hence, this work intrinsically is also limited by those factors given that it was built upon QSAR model. Our exploration of RL is not comprehensive, and we observed that more gains could be achieved through further refining RL reward model.
@@ 4.2 Our models outperform general LLMs
To evaluate PepThink-R1, we considered a series of general-purpose LLMs including both reasoning and non-reasoning models for comparison. Specifically, we include GPT-4o [openai2024gpt4o], DeepSeek-R1-Llama-8B [deepseekai2025deepseekr1incentivizingreasoningcapability], Qwen3-30B-Thinking-2507 [yang2025qwen3technicalreport] and GPT-5 [openai2025gpt5] as LLM baselines. The two GPT models are accessed via API and all other baselines are accessed through open-source weights. The API version of GPT-4o is set to ’gpt-4o-2024-08-06’ and the reasoning effort of GPT-5 is set to minimal. For benchmarking, we adopted vllm[kwon2023efficient] as the inference framework. We used a consistent set of inference hyperparameters. The sampling temperature is set to 0.95, the top-p is set to 0.7 and we restrict the maximum number of generated tokens to 4096.
@@ 4.3 Case study on PepThink-R1 versus PepINVENT
We included PepINVENT, a transformer-based model as a domain-specific baseline and conducted a small case study with four seed peptides to qualitatively compare PepINVENT and PepThink-R1 under challenging settings. Another notably model to compare with is PepTune [tang2025peptune], but we were not able to use it due to limited authorization to for-profit commercial entities. For each seed peptide, we first ranked PepINVENT candidates by the composite total_score derived from LogD, MRT, and SIF, and retained the top-scoring candidate as the PepINVENT examples. For PepThink-R1, we selected one candidate per seed that met the “high” thresholds on all three properties (LogD $\geq$ 4.2, MRT $\geq$ 1.63 h, SIF $t_{1/2}$ $\geq$ 10.1 h). We also include the original seed peptide as the baseline.
@@ 4.4 Interpretable reasoning process by PepThink-R1
Table. 3 shows the exact text generated by PepThink-R1 with visualization of chemical structures of monomers involved in each case. For peptide 1, it mentions replacing glycine with [NH2+][C@@H](Cc1cc(C(C)(C)C)cc(c1Cl)Cl)C(=O)[O-]. The leaving monomer has a matched residue in peptide 1 and the matched residue is replaced by residue of the proposed monomer in peptide generated by PepThink-R1. For peptide 2, the reasoning output instructs replacing leucine with the new monomer which is strictly reflected in the structural change. In both cases, there are more monomer changes proposed by PepThink-R1 that are not captured in the reasoning text. PepINVENT does not have reasoning process, hence, we turned to general reasoning LLMs - GPT-4o and GPT-5 - for comparison. Reasoning outputs from two GPT models for peptide 1 are captured in Table 5. Both outputs from GPT-4o and GPT-5 are generally reasonable high-level strategies for optimizing peptides towards better LogD, MRT, and SIF, and these two outputs share many commonalities, including proposing N-methylation and introducing tert-butyl group. They also roughly cover strategies adopted by PepINVENT and PepThink-R1, but missing detailed monomer structures captured in PepThink-R1 reasoning text. One caveat with GPT reasoning outputs is that they both lack the specific details with the given peptide, which illustrates a limitation of general purpose LLMs and may explain their poor performance (Table.2). The self-consistency discrepancy between great GPT strategy in reasoning output and poor quantitative metrics of resulting peptide SMILES shown in Figure 7 could indicate that either the reasoning process does not reflect the real thinking process of the GPT models, or GPT models are not trained to faithfully carry out their own strategies[chen2023two, lanham2023measuring]. Conversely, PepThink-R1 reasoning output is interpretable, faithfully rooted in the specific case, and also accurately reflected in the final peptide SMILES. Comparing to GPT models, PepThink-R1’s limitation is that its reasoning fails to capture all changes and does not reflect as broad a strategy as GPT is capable of.